Suppose X is a binomial random variable, where n=12 and p = 0.4

compute p < 4

a) 0.5622

b) 0.3453

c) 0.2253

d) 0.4382
Answer #1

Given that ,

p = 0.4

1 - p = 1 - 0.4= 0.6

n = 12

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X < 4) = P(X=0) +P(X=1)+P(X=2)+P(X=3)

= (12 C 0) * 0.40 * (0.6)12+(12 C 1) * 0.41 * (0.6)11+(12 C 2) * 0.42 * (0.6)10+(12 C 3) * 0.43* (0.6)9

=0.2253

probability =0.2253

correct option c